2H NUCLEAR QUADRUPOLE COUPLING OF ORGANIC ACIDS: A COMPARATIVE STUDY

\underline{JENS-UWE GRABOW}, EIKE HENTROP, and DIETER H. SUTTER, Chistian-Albrechts-Universiät, Institut für Physikalische Chemie, Abteilung Chemische Physik, Olshausenstrasse 40, D-24098 Kiel, Germany.

We present an investigation of the deuterium nuclear quadrupole coupling in a series of carbonic acids R-COOH: Among those, the acids with the groups R = CH_{3} and R = CF_{3} are two extreme cases for substituent effects on the charge distribution at the carboxyl group. The deuterium nuclear quadrupole coupling serves as a sensitive probe for the electric field gradient at the location of the acidic proton. The rotational spectra and the quadrupole coupling hyperfine structures of several R-substituted carbonic acids R-COO^{2}H in the range of 6 to 26.5 GHz are reported. Quantum chemical calculations were performed to determine the nuclear quadrupole coupling tensor of ^{2}H. The analysis of the nuclear quadrupole coupling in the rotational spectra provided experimental information on the tensors.